Please help me solve this question by 12p.m. on January 5, 2012:
In an orchard containing 66 trees the number of pear trees is nine more than twice the number of apple trees. How many are there of each kind?

Thanking you in advance,
Helpless student

Answer by mananth(16949)

Apple trees =x
pear trees = 9+2x
total = 66
x+9+2x=66
3x+9=66
-9
3x= 57
/3
x=109
19 apple trees
the balance is pear trees